OpenStudy (anonymous):

I need help please, p^2 + p + 4 = 0

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

im sorry. solve. thats all it gives me. it's in the quadratic section of my book

OpenStudy (maheshmeghwal9):

\[\color{red}{p^2+p+4=0}\]Use quadratic formula; i.e, \[\color{blue}{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}.}\]& remember u must take above quadratic as: -\[\color{green}{ax^2+bx+c=0.}\]Now try to do:)

OpenStudy (maheshmeghwal9):

the two answers depend on \[\huge{\color{purple}{\pm}} \space .\]

OpenStudy (maheshmeghwal9):

take one time plus & one time minus to get two answers:)

OpenStudy (maheshmeghwal9):

Are u trying or nt @raehenson

OpenStudy (anonymous):

i apologize my child was crying

OpenStudy (maheshmeghwal9):

np:)

OpenStudy (anonymous):

i think i understand it, thank you

OpenStudy (maheshmeghwal9):

yw:)